OpenStudy (anonymous):

3x-8.5 = 2.5x-1.25-0.75x Show steps please?

OpenStudy (anonymous):

I got -29. Something tells me that's not right. Help?

OpenStudy (anonymous):

solve by combining like terms first move all terms with variable to one side, then move all constant to the other side then solve for x

OpenStudy (anonymous):

That's what I did. I need someone to check my answer.

OpenStudy (anonymous):

I got -29.

OpenStudy (strawberry17):

I got a different answer. I combined like terms on the right side and then simplified: 3x - 8.5 = 1.75x - 1.25 -1.75x -1.75x ------------------------ 1.25x = 9.75 1.25x 9.75 ----- = ----- 1.25x 1.25 x = ?

OpenStudy (anonymous):

So, you got x=7.8? I've gotten two different answers to this so far. 5.8 and -29. Then you got 7.8 ... I don't know. I must be doing something wrong.

OpenStudy (strawberry17):

yes, I got 7.8 according to how I solved it above. :)

OpenStudy (strawberry17):

how did you get -29?

OpenStudy (anonymous):

I'll show you my steps for that, hold on...

OpenStudy (anonymous):

3x - 8.5 = 2.5x - 1.25 - .75x 3x - 8.5 = 3.25 - 1.25 3x - 3x - 8.5 = 3.25x - 3x -1.25 -8.5 = .25x - 1.25 -8.5 + 1.25 = .25x + 1.25 -7.25 = .25x -7.25/.25 = .25x/.25 x=-29 That's what I got.

OpenStudy (strawberry17):

In the second step, .75x was supposed to be subtracted from 2.5x and not added.. :)

OpenStudy (strawberry17):

do u see that?

OpenStudy (anonymous):

Oh! Dur! I totally blipped that. Let me re-work that and see if I get it right this time...

OpenStudy (strawberry17):

okay! :)

OpenStudy (strawberry17):

Let me know what u get :)

OpenStudy (anonymous):

I got 7.8! I think that's the right answer. Thank you! :D

OpenStudy (strawberry17):

ur welcome! :)

OpenStudy (dwelsher):

1.combine like terms on both sides of the equation(3x-8.5=1.75x-1.25 2.1.25x=7.25--You do know to change the sign when moving to other side of equation? 3.x=7.25/1.25 4. x=5.8 5.If you want to check you work you can substitute 5.8 into the original equation and see that 8.9=8.9

OpenStudy (strawberry17):

sorry, I just checked and @dwelsher is right, it is 5.8. I made a small mistake in my adding... so x = 5.8